# PLRD: Partially Linear Regression Discontinuity Inference

### v1

@@ 5 Formal Results
$\widehat{\tau}_{\texttt{PLRD}}:=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i},\quad\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\underbrace{\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}}_{\text{variance proxy}}+\underbrace{\sup_{\mu_{0},\,\mu_{1}\in\,\mathcal{M}_{\widehat{B}}}\left(\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}}_{\text{worst-case squared bias over}\mathcal{M}_{\widehat{B}}}.$ (8)
If $\widehat{B}$ were a constant chosen a-priori and $\mu^{\prime\prime}_{0}(x)$ were in fact is globally $\widehat{B}$-Lipschitz, we could use existing results for minimax linear regression-discontinuity inference (e.g., Armstrong and Kolesár, 2018; Imbens and Wager, 2019) to directly provide guarantees for $\widehat{\tau}_{\mathtt{PLRD}}$. But here, of course, this is not the case: Our choice of $\widehat{B}$ is learned in step 1 of our algorithm; and Assumption 1 only implies local—not global—Lipschitz continuity in $\mu^{\prime\prime}_{0}(x)$
@@
Suppose that Assumptions 1 and 2 hold and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu^{\prime\prime\prime}(c)|,\varepsilon\}$. Estimate $\widehat{B}$ via local cubic regression: For $w\in\{0,1\}$ run
@@
where $b(\widehat{\gamma})$ is interpreted as the conditional bias of $\widehat{\tau}_{\mathtt{PLRD}}$, and $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$. Furthermore,
@@ 5.1 Proof of Theorem 1
Define $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\sigma}_{i}^{2}$ as in Algorithm 1. It follows from White (1980) that $\widehat{s}(\widehat{\gamma})/s(\widehat{\gamma})\rightarrow_{p}1$; see Lemma 14 for details. Next, fix any $\delta\in(0,1)$. In view of (11) and Slustky’s lemma, $\mathbb{P}(A_{n})\to 1$ where $A_{n}:=\{|b(\widehat{\gamma})|\leq\widehat{b}+\delta\widehat{s}(\widehat{\gamma})\}$. We show in Lemma 13 that for any $s$, $h$, $t$ and $\delta>0$,
@@
where $Z\sim\mathcal{N}(0,1)$. Using this with $s=\widehat{s}(\widehat{\gamma})$, $h=\widehat{h}_{\alpha}$ and $t=\widehat{b}$, we get, for every $n$,
$\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\}\cap A_{n}\right)\geq\inf_{|b|\,\leq\,\widehat{b}}\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha})-\delta\geq 1-\alpha-\delta,$ (12)
@@
$=\liminf_{n\to\infty}\mathbb{P}(\{|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z_{n}|\leq\widehat{h}_{\alpha}\}\cap A_{n})$ (since $\mathbb{P}(A_{n})\to 1$)
$\geq 1-\alpha-\delta.$ (using (12) and Glivenko–Cantelli lemma)
@@ 5.2 Proof of Theorem 3
The optimization problem in Algorithm 1 is inherently related to its convex dual (see Boyd and Vandenberghe (2004) for a definition). We show in Appendix C that the dual problem is given by
@@
In Algorithm 1, we use sample splitting to estimate $B$ (cf. Proposition 2) and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$. However, for ease of exposition, we assume throughout this section that we have $\widehat{B}\rightarrow_{p}B$ and $\widehat{\sigma}^{2}\rightarrow_{p}\sigma^{2}$ independent of the data, and
$\widehat{\theta}_{n}:=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})=\mathop{\rm argmin}_{\theta\,\in\,\Theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B}).$
@@
Although the empirical minimization problem (13) looks similar to a moment-estimation problem, it has some technical complications, e.g., (i) the dual parameter $\theta=(\rho,\lambda)$ involves a function $\rho$ and a vector $\lambda$, (ii) the set $\Theta$ in (14) is not the Cartesian product of a function space and a vector space, and (iii) the population dual problem in (15) involves the sample size $n$. To tackle these issues, we borrow tools from empirical process theory (see, e.g., van der Vaart and Wellner (1996) for a textbook treatment). The following result, proven in Appendix D, characterizes the rate of convergence of the empirical dual optimizer $\widehat{\theta}_{n}$ to the population dual optimizer $\theta_{n}^{*}$; and implies (Corollary 5) that the weights underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy a Lindeberg-type condition.
@@
Let $(\widehat{\rho}_{n},\widehat{\lambda}_{n})=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ be the optimizer of the empirical dual problem (13), and $({\rho}_{n}^{*},\lambda_{n}^{*})=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ be the optimizer of the population dual problem (15). Assume that the pilot bandwidth $\ell$ in Algorithm 1 satisfies (5), and $\widehat{B}\rightarrow_{p}B$. Then, under the conditions of Theorem 1,
@@
Corollary 5 immediately implies that the central limit theorem claimed in Theorem 3 holds; see the proof of Imbens and Wager (2019, Theorem 1) for details. We delegate showing that $s^{2}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-6/7})$ (Lemma 11) and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ (Lemma 12) to Appendix D. Finally, to verify (11), we note that Assumptions 2 and 1 yield $b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})=B\sum_{i=1}^{n}\widehat{\gamma}_{i}\rho(X_{i}-c)$, where
@@
Appendix A Details on Baseline Methods
@@
A.1 Bias-corrected confidence intervals
@@
where $\widehat{V}^{\text{bc}}_{n,p,q}$ is an estimator of the asymptotic variance that accounts for both the variability in estimation and the bias-correction. In Tables 1, 2, 3 and 4, the columns marked as b.c. and rob correspond to the confidence intervals in (18) and (19), respectively.
A.2 Bias-aware confidence intervals
The bias-aware inference approach (e.g., Armstrong and Kolesár, 2018, 2020; Imbens and Wager, 2019) provides a different approach for obtaining confidence intervals for $\tau_{c}$. Note that the LLR estimator takes the form $\widehat{\tau}_{\texttt{LLR}}=\sum_{i=1}^{n}\gamma_{i,\texttt{LLR}}Y_{i}$ for some weights $\gamma_{i,\texttt{LLR}}$ that are functions of the running variable. In fact, the same also holds more generally for the estimator $\widehat{\tau}_{p}(h_{n})$ defined in Section A.1. Next, one makes the observation that although $\widehat{\tau}_{\texttt{LLR}}$ is motivated by a regression problem, the rest of our inference procedure hinges upon on properties that hold for any linear estimators of the form $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$. Given by this observation, we can systematically do better by directly finding weights $\gamma_{i}$ that minimize the mean squared error (MSE) of the estimator $\widehat{\tau}(\gamma)$ uniformly over a class $\mathcal{M}$ of conditional response functions $\mu_{w}$, i.e.,
@@ A.2 Bias-aware confidence intervals
Fixing a value of $M$, we numerically solve the optimization problem (20) and obtain an upper bound on the worst-case bias of the resulting estimator as a by-product. We then follow Imbens and Manski (2004) to construct a bias-aware confidence interval by inflating the Gaussian confidence interval accounting for the worst-case bias of $\tau(\widehat{\gamma})$ under the function class $\mathcal{M}$. These bias-aware confidence intervals provide uniformly valid coverage in finite samples, provided the function class $\mathcal{M}$ and the distribution of $Y_{i}$ given $X_{i}$ is Gaussian (Armstrong and Kolesár, 2018). [^12] If the conditional distribution of $Y_{i}$ is non-Gaussian, one can recover exact finite-sample results using the bias-aware approach by using finite-sample concentration results instead of Gaussian approximation. In practice, however, smoothness constant $M$ is typically unknown; both Armstrong and Kolesár (2020) and Imbens and Wager (2019) suggest using some rule-of-thumb estimates.
A.3 The choice of function class
One of the key insights our paper provides is that the choice of the function class $\mathcal{M}$ in the minimax problem (8) plays a crucial role in determining the conservativeness (and thus the practical usefulness) of the resulting bias-aware method. The function classes used in the existing bias-aware literature (e.g., the classes $\mathcal{M}_{\text{T2},\,M}$ and $\mathcal{M}_{\text{H2},\,M}$ defined in (21) and (22) respectively) are much modest in terms of assumed smoothness as compared to the the bias-corrected methods used in practice (see, e.g., Calonico et al. (2014, Theorem 1)). Here, on the other hand, use the class of all functions $\mu_{w}(\cdot)$ that satisfy Assumption 2 and for which $\mu_{0}^{\prime\prime}(\cdot)$ is $\widehat{B}$-Lipschitz, where we obtain $\widehat{B}$ via a global cubic regression and cross-fitting. While this might appear as a minor modification from a theoretical point of view, it materially affects the length of our intervals in practice, as we demonstrate in Tables 1, 2, 3 and 4.
To illustrate the difference between the conservativeness of our function class versus the function classes used by existing bias-aware approaches, we plot in Fig. 3 the regression functions corresponding to the worst-case bias obtained by our method and $\mathcal{M}_{\text{H2},\,M}$ as used in RDHonest. This plot demonstrates that the classes $\mathcal{M}_{\text{H2},\,M}$ can often be unreasonably broad, and include regression functions that are ex ante implausible or rarely occur in practice. This explains why the resulting bias-aware confidence intervals can be substantially more conservative than the bias-corrected ones (see, e.g.Table 3). The empirical success of our PLRD method that we demonstrate in Section 4 is, in fact, rooted in this choice of the function class over which one computes the worst-case MSE.
Figure 3: Functions $\mu_{w}(\cdot)$ with worst-case bias in the RDHonest and PLRD function classes. The top two plots use the $m_{\text{quad}}$ example from Imbens and Kalyanaraman (2012), while the bottom two plots use the data from Lee (2008). In each plot, black lines show $\mu_{w}(\cdot)$ functions with maximum (solid) or minimum (dotted) bias in the respective function class. This figure illustrates how the RDHonest function class can be too modest for many regression discontinuity designs in practice.
A.4 Ablation analysis

### v2

@@ 5 Formal Results
$\widehat{\tau}_{\texttt{PLRD}}:=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i},\quad\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\underbrace{\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}}_{\text{variance proxy}}+\underbrace{\sup_{\mu_{0},\,\mu_{1}\in\,\mathcal{M}_{\widehat{B}}}\left(\sum_{i=1}^{n}{\gamma}_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}}_{\text{worst-case squared bias over }\mathcal{M}_{\widehat{B}}}$. (9)
If $\widehat{B}$ were a constant chosen a-priori and $\mu^{\prime\prime}_{0}(x)$ were in fact is globally $\widehat{B}-$Lipschitz, we could use existing results for minimax linear regression-discontinuity inference (e.g., Armstrong and Kolesár, 2018; Imbens and Wager, 2019) to directly provide guarantees for $\widehat{\tau}_{\mathtt{PLRD}}$. But here, of course, this is not the case: Our choice of $\widehat{B}$ is learned in step 1 of our algorithm, and Assumption 1 only implies local—not global—Lipschitz continuity in $\mu^{\prime\prime}_{0}(x)$ .
@@
Suppose that Assumptions 1 and 2 hold and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu_{0}^{\prime\prime\prime}(c)|,|\mu_{1}^{\prime\prime\prime}(c)|,\varepsilon\}$. Run the local cubic regression:
$\widehat{\beta}:=\mathop{\rm argmin}_{\beta\in\mathbb{R}^{6}}\sum_{\begin{subarray}{c}|X_{i}-c|\leq\ell\end{subarray}}\left(Y_{i}-\beta_{W_{i},0}-\beta_{W_{i},1}(X_{i}-c)-\frac{\beta_{2}}{2}(X_{i}-c)^{2}-\frac{\beta_{3}}{6}(X_{i}-c)^{3}\right)^{2},$ (10)
with bandwidth $\ell$ scaling as in (6), and set [^15] As an additional safe-guard against small sample sizes, we calculate a standard error estimate $\mathrm{s.e.}(\widehat{\beta}_{3})$ and use the conservative yet asymptotically consistent estimate $\widehat{B}=\max\{|\widehat{\beta}_{3}\pm 1.96\cdot\mathrm{s.e.}(\widehat{\beta}_{3})|,\varepsilon\}$ in our software. $\widehat{B}=\max\{|\beta_{3}|,\varepsilon\}$. Then, $\widehat{B}\rightarrow_{p}B$ as $n\to\infty$.
@@
$\left(\widehat{\tau}_{\mathtt{PLRD}}-b(\widehat{\gamma})-\tau_{c}\right)\,\big/\,{s}(\widehat{\gamma})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,\,1),\ \ \ \ {s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2},\ \ \ \ b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\,\mu_{W_{i}}(X_{i})-\tau_{c},$ (11)
@@ 5.1 Proof of Theorem 1
[…]
@@
[…]
$\begin{split}\text{$\mathbf{1}$}_{A_{n}}\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\,\Big|\,(X_{i},Y_{i})_{i=1}^{n}\right)&\geq\text{$\mathbf{1}$}_{A_{n}}\inf_{|b|\,\leq\,\widehat{b}\,+\,\delta\widehat{s}(\widehat{\gamma})}\mathbb{P}\left(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\right)\\ &\geq\text{$\mathbf{1}$}_{A_{n}}\left(\inf_{|b|\,\leq\,\widehat{b}}\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha})-\delta\right)\\ &\geq\text{$\mathbf{1}$}_{A_{n}}(1-\alpha-\delta),\end{split}$ (13)
@@
$=\liminf_{n\to\infty}\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\right)$
$\geq\liminf_{n\to\infty}\mathbb{P}(\{|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\}\cap A_{n})$
$\geq 1-\alpha-\delta.$ (using (13) and $\mathbb{P}(A_{n})\to 1$)
@@ 5.2 Proof of Theorem 3
The optimization problem in Algorithm 1 is inherently related to its convex dual (see Boyd and Vandenberghe (2004) for a definition). We show in Appendix D that the dual problem is given by
@@
In Algorithm 1, we use sample splitting to estimate $B$ (cf. Proposition 2) and the conditional variance $\sigma^{2}=\mathbb{E}[{\rm Var}(Y_{i}\mid X_{i})]$. However, for ease of exposition, we assume throughout this section that we have $\widehat{B}\rightarrow_{p}B$ and $\widehat{\sigma}^{2}\rightarrow_{p}\sigma^{2}$ independently of the data, and
[…]
@@
Although the empirical minimization problem (14) looks similar to a moment-estimation problem, it has some technical complications, e.g., (i) the dual parameter $\theta=(\rho,\lambda)$ involves a function $\rho$ and a vector $\lambda$, (ii) the set $\Theta$ in (15) is not the Cartesian product of a function space and a vector space, and (iii) the population dual problem in (16) involves the sample size $n$. To tackle these issues, we borrow tools from empirical process theory (see, e.g., van der Vaart and Wellner (1996) for a textbook treatment). The following result, proven in Appendix E, characterizes the rate of convergence of the empirical dual optimizer $\widehat{\theta}_{n}$ to the population dual optimizer $\theta_{n}^{*}$; and implies (Corollary 5) that the weights underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy a Lindeberg-type condition.
@@
$\|\widehat{\rho}_{n}-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}^{2}=\mathcal{O}_{p}\left(\frac{\ell^{\,6}}{n^{1/4+6/7}}+\frac{\ell^{\,3}}{n^{1+3/7}}\right).$
@@
Corollary 5 immediately implies that the central limit theorem claimed in Theorem 3 holds; see the proof of Imbens and Wager (2019, Theorem 1) for details. We delegate showing that $s^{2}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-6/7})$ (Lemma 12) and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ (Lemma 13) to Appendix E. Finally, to verify (12), we note that Assumptions 2 and 1 yield $b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})=B\sum_{i=1}^{n}\widehat{\gamma}_{i}\rho(X_{i}-c)$, where
@@
Appendix A Safeguard against Curvature Changes
While Assumption 2 provides a parsimonious and empirically realistic model for treatment effect heterogeneity in many applications, it can in principle be violated. To safeguard against such violations, our software incorporates an automatic diagnostic that tests for difference in curvatures. The diagnostic works as follows. Before running the main PLRD procedure, we conduct an ANOVA test comparing a restricted cubic polynomial model (where treatment interacts only linearly with $X_{i}-c$, consistent with Assumption 2) against an unrestricted model (where treatment interacts with all polynomial terms in $X_{i}-c$). If the test rejects at significance level $\alpha^{\prime}$ (default 0.1%), the method proceeds with a more flexible specification that only assumes Assumption 1 but not Assumption 2. Otherwise, it uses the default specification under both assumptions. This adaptive approach helps explain the strong empirical performance of plrd across all simulation settings in Table 3, including those where Assumption 2 fails.
When the flexible specification is selected, estimation of the Lipschitz constant $B$ changes to accommodate potentially different curvature on either side of the threshold. Rather than fitting a single cubic polynomial as in (10), we fit separate cubic regressions for treated and control units:
Proposition 6 (Consistent estimation of the Lipschitz constant $B$ under different curvatures).
Suppose that Assumption 1 holds and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu_{0}^{\prime\prime\prime}(c)|,|\mu_{1}^{\prime\prime\prime}(c)|,\varepsilon\}$. Estimate $\widehat{B}$ via local cubic regression: For $w\in\{0,1\}$ run
With this estimated $\widehat{B}$ in hand, the subsequent minimax optimization in Algorithm 1 adjusts the constraint set to allow for different curvature functions $\rho_{0}$ and $\rho_{1}$ on either side of the cutoff (see (4) versus (3)). The optimization problem remains computationally tractable via quadratic programming, and the remainder of the PLRD procedure—including cross-fitting, variance estimation, and confidence interval construction—proceeds similarly to the default case. We note that a formal proof of asymptotic validity under this flexible specification would follow similar arguments to those in Theorem 1, with nearly identical technical steps that are omitted here.
Appendix B Details on Baseline Methods
@@
B.1 Bias-Corrected Confidence Intervals
@@
where $\widehat{V}^{\text{bc}}_{n,p,q}$ is an estimator of the asymptotic variance that accounts for both the variability in estimation and the bias-correction. In Tables 1, 2 and 3, the columns marked as rdrobut correspond to the confidence intervals computed using (21).
B.2 Bias-Aware Confidence Intervals
The bias-aware inference approach (e.g., Armstrong and Kolesár, 2018, 2020; Imbens and Wager, 2019) provides a different approach for obtaining confidence intervals for $\tau_{c}$. Note that the LLR estimator takes the form $\widehat{\tau}_{\texttt{LLR}}=\sum_{i=1}^{n}\gamma_{i,\texttt{LLR}}Y_{i}$ for some weights $\gamma_{i,\texttt{LLR}}$ that are functions of the running variable. In fact, the same also holds more generally for the estimator $\widehat{\tau}_{p}(h_{n})$ defined in Section B.1. Next, one makes the observation that although $\widehat{\tau}_{\texttt{LLR}}$ is motivated by a regression problem, the rest of our inference procedure hinges upon on properties that hold for any linear estimators of the form $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$. Given by this observation, we can systematically do better by directly finding weights $\gamma_{i}$ that minimize the mean squared error (MSE) of the estimator $\widehat{\tau}(\gamma)$ uniformly over a class $\mathcal{M}$ of conditional response functions $\mu_{w}$, i.e.,
@@ B.2 Bias-Aware Confidence Intervals
[…]
B.3 The Choice of Function Class
One of the key insights our paper provides is that the choice of the function class $\mathcal{M}$ in the minimax problem (9) plays a crucial role in determining the conservativeness (and thus the practical usefulness) of the resulting bias-aware method. The function classes used in the existing bias-aware literature (e.g., the classes $\mathcal{M}_{\text{T2},\,M}$ and $\mathcal{M}_{\text{H2},\,M}$ defined in (23) and (24) respectively) are much modest in terms of assumed smoothness as compared to the the bias-corrected methods used in practice (see, e.g., Calonico et al. (2014, Theorem 1)). Here, on the other hand, use the class of all functions $\mu_{w}(\cdot)$ that satisfy Assumption 2 and for which $\mu_{0}^{\prime\prime}(\cdot)$ is $\widehat{B}-$Lipschitz, where we obtain $\widehat{B}$ via a global cubic regression and cross-fitting. While this might appear as a minor modification from a theoretical point of view, it materially affects the width of our intervals in practice, as we demonstrate in Tables 1, 2 and 3.
To illustrate the difference between the conservativeness of our function class versus the function classes used by existing bias-aware approaches, we plot in Figure 2 the regression functions corresponding to the worst-case bias obtained by our method and $\mathcal{M}_{\text{H2},\,M}$ as used in RDHonest. This plot demonstrates that the classes $\mathcal{M}_{\text{H2},\,M}$ can often be unreasonably broad, and include regression functions that are ex ante implausible or rarely occur in practice. This explains why the resulting bias-aware confidence intervals can be substantially more conservative than the bias-corrected ones (see, e.g.Tables 1, 2 and 3). The empirical success of our PLRD method that we demonstrate in Section 4 is, in fact, rooted in this choice of the function class over which one computes the worst-case MSE.
Figure 2: Functions $\mu_{w}(\cdot)$ with worst-case bias in the RDHonest and PLRD function classes. The top two plots use the $m_{\text{quad}}$ example from Imbens and Kalyanaraman (2012), while the bottom two plots use the data from Lee (2008). In each plot, black lines show $\mu_{w}(\cdot)$ functions with maximum (solid) or minimum (dotted) bias in the respective function class. This figure illustrates how the RDHonest function class may accommodate in-hindsight-unreasonable functions for several practical settings.
B.4 Ablation Analysis